Find Mean and Variance, if they exist in the following distribution:

fx={6x1-x   ,0<x<1 0  ,              elsewhere 


1
Expert's answer
2021-07-07T14:03:45-0400
"f(x) = \begin{cases}
 6x(1-x) & 0<x<1 \\
 0 & elsewhere
\end{cases}"

"E(X)=\displaystyle\int_{-\infin}^{\infin}xf(x)dx=\displaystyle\int_{0}^{1}6x^2(1-x)dx"

"=[2x^3-1.5x^4]\begin{matrix}
 1 \\
 0
\end{matrix}=0.5"

"E(X^2)=\displaystyle\int_{-\infin}^{\infin}x^2f(x)dx=\displaystyle\int_{0}^{1}6x^3(1-x)dx"

"=[1.5x^4-1.2x^5]\begin{matrix}
 1 \\
 0
\end{matrix}=0.3"

"Var(X)=E(X^2)-(E(X))^2=0.3-(0.5)^2=0.05"